In 1987, there were 4.946 million people registered
to vote in a state. Of these, x million
were registered democrats, y million were registered
republicans, and z million were registered
as independent. The value of x is 0.599
million more than the value of y. The value
of y is 1.356 million more than z. How many
democrats were registered?

Answer by TimothyLamb(4379)

x + y + z = 4.946
x = y + 0.599
y = z + 1.356
---
put the system of linear equations into standard form
---
x + y + z = 4.946
x - y = 0.599
y - z = 1.356
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = democrats = 2.5 million
y = republicans = 1.901 million
z = independent = 0.545 million
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---